Could someone help me understand this problem?
Determine the number of solutions and classify the type of solutions for the following equation and justify the answer.
sqrt2 x^2 - 4x - 7 sqrt2=0

Answer by RAY100(1637)

(sqrt2)x^2 -4x -7sqrt2 =0
.
1.414x^2 -4x -9.899 =0
.
Discriminant(D) = b^2 -4ac,,,,where a=1.414,,b= -4,,,c=-9.899
.
D= { (-4)^2 -4(1.414) (-9.899) } = 72
.
Rules for Discriminant are:
.
D=0,,,one solution, curve intersects x axis in one location
.
D>0,,, two solutions,,curve intersects x axis in two locations
.
D<0,,,no Real solutions,,,curve does not intersect x axis
.
Finishing the quadratic eqn finds ,,,x= -1.586, +4.415